Algebra

An article is marked at ₹ 940. If it is sold for ₹ 799, then find the discount per cent. *

Tanishqa Nambiar
5 Years agoGrade

3 Answers

Tanishqa Nambiar
5 Years ago
cost price is = 940 rupees 
sold for=799 rupees
find discount percent
to find discount amount=940-799=141
so formula is = discount amount divided by marked price(base price) x 100
which means 141 by 940(base price) x 100= 15 percentage
Tanishqa Nambiar
5 Years ago
The salary of Manish is Rs 10000. His salary gets increased by 10%. Find his increased salary. 
old salary=10000 rupees 
new salary= x
salary is increased by =10 percent
10percent of 10000=10 by 100 x 10000=1000
salary is increased by 10% .
so, present salary = 10,000 + 1000 = 11,000
Ram Kushwah
5 Years ago
Here Base price = Marked price
Now we knw that
Discount=Marked price – Selling price
=Rs 940-799
=Rs 141
 Now Discount % =(Discount) / ( Marked price) x 100
=(141/940) x 100
=0.15 x 100
=15 %
Thus the discount = 15 %